OpenStudy (anonymous):

LCD OF 9over10 and 5over8 ?

OpenStudy (anonymous):

9/10=5/8 Since 10 & 8 have a common multiple yet 9 & 5 do not, multiply 10*8 = 80. The LCD would then be 80. 72/80=50/80